Operations Report – January 25th

Hi Mission Support!

Crew 203 Operations Report 25-JAN-2019

SOL: 13

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #11 and suit #6 batteries don’t charge properly.

Notes on non-nominal systems: None

Generator (hours run): 0hr 0min; Not turned on last night (24Jan2019); Turned On today (25Jan2019) at 16:45.

Solar SOC – Turned on (24Jan2019) 86%; Turned off (25Jan2019) 100%; Turned on (25Jan2019) 91%

Diesel Reading – Under 1/2 tank.

Propane Reading – 80%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 40%; 215 gallons (Refilled)

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 33%; 98 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 20:20 to refill tank

Water Meter: 01404805

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used.

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Not used.

Hours: 46.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used.

Hours: 72.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Used to get water from Hanksville.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #10 and Suit #6 batteries don´t charge.

Summary of Hab operations: RV Digest liquid is almost empty. Water tank was refilled with 125 gallons, it was refilled to prevent a damage on the submersible pump and to get a shower before leaving Mars.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

Freddy